# DESY 24-073 May 2024 CERN-TH-2024-068 ZU-TH 26/24

### v1

@@ IV Resummation
The relevant physical settings and cuts on $Q\equiv m_{\ell\ell}$ and $Y\equiv Y_{\ell\ell}$ are given by the input file in App. D. Note that we always cut on the lepton pseudorapidities $|\eta_{\ell_{1},\ell_{2}}|\leq 2.5$ in addition. The cuts on the lepton transverse momenta are given in Eqs. (3) and (12) for staggered and product cuts, respectively.
@@ IV.2 Numerical results for the non-singular cross section
In Fig. 10 we show the fixed-order perturbative coefficients of the non-singular cross section including the respective factors of $\alpha_{s}^{n}$ in the case of the staggered cuts in Eq. (3). Within the numerical uncertainty, which is indicated by the blue band and driven by the SCETlib beam function interpolation uncertainty, these are all compatible with an asymptotic behavior of two straight lines approaching a constant as $\Delta p_{T}\to 0$ with different slopes. The LO non-singular cross section is compatible with zero within a relative accuracy of $10^{-5}$, validating the SCETlib physics inputs against those used at fixed order (in NNLOJET here).
For the case of product cuts defined in Eq. (12) the non-singular cross sections at LO, NLO and NNLO are plotted in Fig. 11. The only notable feature compared to Fig. 10 is a faster (possibly quadratic) rise of the $\alpha_{s}^{2}$ coefficient towards large positive values of $\Delta p_{T}$, but this is at the scale of a few tenths of a picobarn.
@@ IV.3 Matched predictions
The resummation is under good theoretical control, as evidenced by the matching uncertainty $\Delta_{\mathrm{match}}$ which we plot together with the matched predictions. Here $\Delta_{\mathrm{match}}$ is estimated by profile scale variations exactly as described in Ref. Ebert:2020dfc and features excellent perturbative convergence and coverage. As an important cross check, we have verified that our resummation and matching setup lead to a net resummation effect compatible with zero when applied to the inclusive cross section within the ATLAS $Q^{2}$ and $Y$ bins. Explicitly, we find
@@
where ±$\pm_{\mathrm{match}}$ indicates the matching uncertainty $\Delta_{\mathrm{match}}$ at the given order. We note that the profile scale variations are not currently tuned to preserve the total inclusive cross section as computed at the reference value of $\mu_{R}=\mu_{F}$, which is reflected in the non-zero value of $\Delta_{\mathrm{match}}$ above. These variations are therefore still somewhat redundant with standard $(\mu_{R},\mu_{F})$ variations and would have to be constrained through the total inclusive cross section in an improved analysis.
@@
Another uncertainty intrinsic to the resummed prediction is given by the impact of non-perturbative transverse momentum dependent (TMD) physics and the Landau pole in the inverse Fourier transform, which is estimated by variations of a cutoff parameter $\Lambda_{\mathrm{fr}}$ defined in Ref. Ebert:2020dfc . We find that the resulting $\Delta_{\mathrm{NP}}$ is not even resolved at the current relative numerical uncertainty, which is ${\cal O}(10^{-3})$. This is expected because for typical $q_{T}\sim 25\,\text{GeV}$ in the baseline symmetric cuts (which dominate the total fiducial cross section here), the effect of non-perturbative TMD physics is suppressed by $(\Lambda_{\mathrm{QCD}}/q_{T})^{2}\sim(500\,\text{MeV}/25\,\text{GeV})^{2}\sim 4\times 10^{-4}$.
@@
Fig. 13 shows our results for the matched cross sections for the case of the product cuts defined in Eq. (12), also compared to the NNLOJET fixed-order results. They are again almost identical to their counterparts for asymmetric cuts in Fig. 12. A few comments on the product cuts are in order. Ref. Salam:2021tbm focused on constructing $1\to 2$ fiducial cuts that would be free of $\mathcal{O}(q_{T})$ linear power corrections at the level of the differential spectrum as $q_{T}\to 0$, which were found to reduce ambiguities in the perturbative series for the case of Higgs production with $\Delta p_{T}\sim 10$ GeV. Furthermore, the absence of linear power corrections in the acceptance is beneficial for mitigating bias in fixed-order subtraction methods. However, the sensitivity of the total fiducial cross section to resummation effects remains present through terms of the form $x\ln^{n}x$ for $x=\Delta p_{T}/M_{Z}\ll 1$. This is evident from our numerical results in Fig. 8, where the unphysical behavior for $\Delta p_{T}\rightarrow 0$ of the fixed-order cross section persists also for product cuts. We note that this observation does not straightforwardly follow from the definition of product cuts. Although they trivially coincide with the symmetric cuts for $\Delta p_{T}\to 0$, they might have displayed a different asymptotic behavior approaching that limit. In the language of Ref. Ebert:2020dfc , there are two classes of fiducial power corrections, either “linear” or “leptonic”, which do not have to coincide. “Leptonic” power corrections in $q_{T}/\Delta p_{T}=xq_{T}/M_{Z}$ are in general present and lead to sensitivity to low values of $q_{T}$ at any power in $q_{T}/\Delta p_{T}$. Ultimately, the only way to cure the pathological behavior in the $\Delta p_{T}=0$ limit is to resort to resummation, due to the presence of $x\ln x$ power corrections for $q_{T}/M_{Z}\lesssim x$, as already observed in Frixione:1997ks .
@@ VI Conclusions and Discussion
We also considered product cuts Salam:2021tbm , which replace the separate cuts on the leading and sub-leading lepton transverse momenta with cuts on their product and the sub-leading lepton instead. For values of $\Delta p_{T}\sim 10$ GeV these cuts mitigate the ambiguities of the perturbative series present with (a)symmetric cuts, which can be particularly relevant for processes with Casimir enhancement, notably Higgs production Billis:2021ecs . However, we find that their introduction does not address the pathological behavior for $\Delta p_{T}\rightarrow 0$, which can only be addressed by resummation. We found that resummation indeed removes this effect, although a kink in the cross section is still present for $\Delta p_{T}=0$, essentially due to the way we defined the observable for this study. At NNLO the differences induced by resummation are small for all the values of $\Delta p_{T}$ considered, ranging from 00 to a few GeV, shifting the fixed-order result by sub-percent values that are well below the current experimental uncertainties. It remains to be seen whether such conclusion still holds beyond NNLO, although recent studies Salam:2021tbm ; Chen:2022cgv indicate that the effect of resummation of linear power corrections remains moderate also at N3LO. Given the excellent job done by the available fixed-order codes in reproducing the results of resummation and the substantial effort needed to implement the relevant modifications when modeling signal and background in experimental analyses, we foresee only a marginal improvement in adopting a different set of cuts in the Drell-Yan case for the measurement of fiducial cross sections.

### v2

@@ IV Resummation
The relevant physical settings and cuts on $Q\equiv m_{\ell\ell}$ and $Y\equiv Y_{\ell\ell}$ are given by the input file in App. D. Note that we always cut on the lepton pseudorapidities $|\eta_{\ell_{1},\ell_{2}}|\leq 2.5$ in addition. The cuts on the lepton transverse momenta are given in Eqs. (3) and (12) for linear asymmetric and product cuts, respectively.
@@ IV.2 Numerical results for the non-singular cross section
In Fig. 13 we show the fixed-order perturbative coefficients of the non-singular cross section including the respective factors of $\alpha_{s}^{n}$ in the case of the linear asymmetric cuts in Eq. (3). Within the numerical uncertainty, which is indicated by the blue band and driven by the SCETlib beam function interpolation uncertainty, these are all compatible with an asymptotic behavior of two straight lines approaching a constant as $\Delta p_{T}\to 0$ with different slopes. The LO non-singular cross section is compatible with zero within a relative accuracy of $10^{-5}$, validating the SCETlib physics inputs against those used at fixed order (in NNLOJET here). We stress that the nonsingular cross section in Eq. (14) is independent of the resummation formalism used. This is because the singular spectrum, whose integral is subtracted from the total fixed-order cross section in Eq. (14), is specified by evaluating the underlying hadronic structure functions in fixed-order perturbation theory and expanding them to leading power in $q_{T}/Q$, while keeping the exact dependence on $q_{T}$ in the leptonic phase space integral. The choice to evaluate the structure functions in the Collins-Soper frame and the choice to expand in $1/Q$ (instead of the inverse of the transverse mass) amount to quadratic power corrections. Due to their lower degree of divergence at the level of the $q_{T}$ spectrum, these cannot induce logarithmic dependence on $\Delta p_{T}$, and thus these choices (like the choice of $q_{T}^{\mathrm{off}}$) only affect the constant terms and linear slopes in Eq. (14). For the case of product cuts defined in Eq. (12) the non-singular cross sections at LO, NLO and NNLO are plotted in Fig. 14. The only notable feature compared to Fig. 13 is a faster (possibly quadratic) rise of the $\alpha_{s}^{2}$ coefficient towards large positive values of $\Delta p_{T}$, but this is at the scale of a few tenths of a picobarn.
@@ IV.3 Matched predictions
The resummation is under good theoretical control, as evidenced by the matching uncertainty $\Delta_{\mathrm{match}}$, which we plot together with the matched predictions in the left column of Fig. 15. Here $\Delta_{\mathrm{match}}$ is estimated by profile scale variations using the formalism and default settings introduced in Ref. Ebert:2020dfc , and features excellent perturbative convergence and coverage. For later reference, we briefly describe the salient features of this default matching formalism. Ref. Ebert:2020dfc employs a hybrid profile scale technique Lustermans:2019plv , where up to a value of $q_{T}=x_{1}Q$ the resummation is fully turned on. In this first region, the spectrum is obtained from a numerical Fourier transform of the resummed $b_{T}$-space cross section. The beam and soft functions are evaluated at the so-called canonical boundary values for their virtuality and rapidity scales,
$\mu_{S}(b_{T})=\mu_{B}(b_{T})=\nu_{S}(b_{T})=\frac{b_{0}}{b_{T}}\,,\qquad\nu_{B}=Q\,,$
where $b_{0}=2e^{-\gamma_{E}}$ and $\gamma_{E}$ is the Euler-Mascheroni constant. An identical result for the resummed $b_{T}$-space cross section is obtained when choosing canonical $\mu$ and Collins-Soper scales $\zeta$ in TMD factorization. For $q_{T}\geq x_{1}Q$, the resummation is slowly turned off by promoting the scales in Eq. (IV.3) to be functions of $q_{T}$ that continuously and monotonically approach the fixed-order value $\mu_{R}=\mu_{F}=Q$ for all values of $b_{T}$ as $q_{T}\to Q$. Specifically, one picks another point $q_{T}=x_{3}Q>x_{1}Q$ beyond which all scales are equal to $Q$, the resummation is fully off, and the fixed-order result for the spectrum is recovered exactly. A third intermediate point $q_{T}=x_{2}Q$ with $x_{1}<x_{2}<x_{3}$ relates to the precise functional form of the transition and governs how quickly it departs from the canonical resummation region. The default settings for these parameters from Ref. Ebert:2020dfc , which were determined quantitatively by comparing the size of leading-power singular and the nonsingular terms in the $q_{T}$ spectrum, are given by
$(x_{1},x_{2},x_{3})=(0.3,0.6,0.9)\,.$ (19)
The matching uncertainty $\Delta_{\mathrm{match}}$ is then estimated by varying them as
$(x_{1},x_{2},x_{3})\in V_{\mathrm{match}}=\{(0.4,0.75,1.1),\,(0.2,0.45,0.7),\,(0.4,0.55,0.7),\,(0.2,0.65,1.1)\}\,,$ (20)
i.e., one either shifts the transition region up, shifts it down, condenses it, or stretches it out. Notably, the relative size of singular and nonsingular contributions from which these default choices were determined was assessed at the level of the $q_{T}$ spectrum in Ref. Ebert:2020dfc in order to obtain an optimal prediction for its shape. By contrast, neither the central value of the $x_{i}$ nor their variations were tuned to exactly preserve the fixed-order prediction for the integral of the inclusive $q_{T}$ spectrum, i.e., the total inclusive cross section. As an important cross check, we have therefore explicitly verified that the above default resummation and matching setup nevertheless leads to a net resummation effect compatible with zero when applied to the inclusive cross section within the ATLAS $Q^{2}$ and $Y$ bins. Explicitly, we find
@@
where ±$\pm_{\mathrm{match}}$ indicates the matching uncertainty $\Delta_{\mathrm{match}}$ at the given order.
As an alternative to the default spectrum-level setup from Ref. Ebert:2020dfc , one may instead consider profile scales that are chosen such that the matching exactly preserves the total inclusive cross section as computed at a given fixed order in perturbation theory. We stress that because the default matching is implemented through profile scales, the nonzero resummation effects in Eq. (IV.3) are by construction also beyond the order at which the fixed-order calculation is truncated. To do so we maintain the original functional form of the transition as given in Ref. Ebert:2020dfc , but choose the transition points $x_{i}$ under the constraint that for the central prediction (which uses Eq. (IV.3) in the canonical region $q_{T}\leq x_{1}Q$), the integral of the matched cross section up to $q_{T}=x_{3}Q$ exactly recovers the fixed-order value. We further choose to hold $x_{1}=0.3$ fixed for the central prediction such that the size of the canonical region coincides with the default settings. By manually scanning the dependence on the other two parameters, we then find that
$(x_{1},x_{2},x_{3})_{\mathrm{constr.}}=(0.3,0.9,1.2)$ (22)
exactly recovers (for the settings used in this paper, and in particular the PDF set at hand) the NNLO total inclusive cross section from the integral of the N3LL$+$NNLO inclusive $q_{T}$ spectrum.
It is interesting to ask how an uncertainty estimate based on scale variations can still be obtained for the $q_{T}$ spectrum in this case, as well as for derived quantities like fiducial cross sections that feature sensitivity to resummation effects, since the matching uncertainty has largely become trivial by imposing the above constraint. In particular, these additional variations should probe possible changes of the shape of the spectrum in the canonical region itself, which can also serve as a proxy for the difference between different resummation formalisms. To test this we adapt another, separate component of the total uncertainty estimate for the $q_{T}$ spectrum from Ref. Ebert:2020dfc , specifically, the so-called resummation uncertainty $\Delta_{\mathrm{res}}$. Conventionally, $\Delta_{\mathrm{res}}$ is estimated by performing a large set of variations where the four canonical scales in Eq. (IV.3) are varied independently or jointly by factors of $2$ around their central values Stewart:2013faa . Taking an envelope of these variations then provides one of the most detailed estimates of the residual uncertainty on the spectrum in resummed perturbation theory available within the scale variation paradigm. While these variations are smoothly turned off as $q_{T}$ increases such that the fixed-order prediction in the tail of the spectrum is unaffected by them, they again do not necessarily preserve the total integral. While they probe variations from the canonical Sudakov shape in depth, they are not necessarily a pure shape variation of the spectrum. To complete our uncertainty estimate in the case where the $x_{i}$ are constrained, we also subject the variations entering $\Delta_{\mathrm{res}}$ to the integral constraint by adjusting the $x_{i}$ for each variation. Restricting to a subset of representative variations to keep the complexity of the problem manageable, we find
$V_{\mathrm{res}}^{\rm constr.}$ =$=\bigl{\{}(\mu_{B}^{\mathrm{down}},\mu_{S}^{\mathrm{down}},0.55,1.0,1.2),\,(\mu_{B}^{\mathrm{up}},\mu_{S}^{\mathrm{up}},0.25,0.65,1.2),\,$
$\qquad(\nu_{B}^{\mathrm{down}},0.35,0.9,1.2),\,(\nu_{B}^{\mathrm{up}},0.4,0.8,1.2)\bigr{\}}\,,$ (23)
where $\mu_{X}^{\mathrm{up,down}}$ indicates a soft or beam function scale variation as described in Ref. Ebert:2020dfc , and the three values quoted are the respective $x_{i}$. Note that in this case, we have held $x_{3}$ fixed at the central value determined earlier to ensure that all variations in $V_{\mathrm{res}}^{\rm constr.}$ collapse onto the fixed-order spectrum at the same point, and instead compensated the changes by mainly varying $x_{1}$, i.e., the size of the purely canonical region.
The results of the above exercise are shown in the right column of Fig. 15. Compared to the left column, the central values of the matched predictions are shifted up, largely compensating – now at the fiducial level – the small offsets we reported in Eq. (IV.3). Note that since we are mainly interested in the behavior at the highest order in perturbation theory, we have used the same constrained variations from above at all three orders shown in Fig. 15 for expediency. This is the reason for the much larger uncertainty bands at lower orders. Nevertheless, as can be seen from the bottom right panel, we still find a small net resummation effect on the central value at N3LL$+$NNLO that is mainly compatible with zero, but just outside the refined uncertainty estimate for some values of $\Delta p_{T}$. While a complete study of this effect at the next higher order is beyond the scope of this paper, we anticipate that the effect will become more significant at higher orders since the estimate for $\Delta_{\mathrm{res}}^{\mathrm{constr.}}$ will be reduced as the residual scale dependence decreases in the resummed cross section, while the baseline fixed-order prediction picks up two additional logarithms of $\Delta p_{T}/Q$ at each order according to Eq. (IV.1).
@@
Another uncertainty intrinsic to the resummed prediction is given by the impact of non-perturbative transverse momentum dependent (TMD) physics and the Landau pole in the inverse Fourier transform, which is estimated by variations of a cutoff parameter $\Lambda_{\mathrm{fr}}$ defined in Ref. Ebert:2020dfc . We find that the resulting $\Delta_{\mathrm{NP}}$ is not even resolved at the current relative numerical uncertainty, which is ${\cal O}(10^{-4})$. This is expected because for typical $q_{T}\sim 25\,\text{GeV}$ in the baseline symmetric cuts (which dominate the total fiducial cross section here), the effect of non-perturbative TMD physics is suppressed by $(\Lambda_{\mathrm{QCD}}/q_{T})^{2}\sim(500\,\text{MeV}/25\,\text{GeV})^{2}\sim 4\times 10^{-4}$.
@@
Fig. 16 shows our results for the matched cross sections for the case of the product cuts defined in Eq. (12), also compared to the NNLOJET fixed-order results. They are again almost identical to their counterparts for linear asymmetric cuts in Fig. 15. A few comments on the product cuts are in order. Ref. Salam:2021tbm focused on constructing $1\to 2$ fiducial cuts that would be free of $\mathcal{O}(q_{T})$ linear power corrections at the level of the differential spectrum as $q_{T}\to 0$, which were found to reduce ambiguities in the perturbative series for the case of Higgs production with $\Delta p_{T}\sim 10$ GeV. Furthermore, the absence of linear power corrections in the acceptance is beneficial for mitigating bias in fixed-order subtraction methods. However, the sensitivity of the total fiducial cross section to resummation effects remains present through terms of the form $x\ln^{n}x$ for $x=\Delta p_{T}/M_{Z}\ll 1$. This is evident from our numerical results in Fig. 8, where the unphysical behavior for $\Delta p_{T}\rightarrow 0$ of the fixed-order cross section persists also for product cuts. We note that this observation does not straightforwardly follow from the definition of product cuts. Although they trivially coincide with the symmetric cuts for $\Delta p_{T}\to 0$, they might have displayed a different asymptotic behavior approaching that limit. In the language of Ref. Ebert:2020dfc , there are two classes of fiducial power corrections, either “linear” or “leptonic”, which do not have to coincide. “Leptonic” power corrections in $q_{T}/\Delta p_{T}=xq_{T}/M_{Z}$ are in general present and lead to sensitivity to low values of $q_{T}$ at any power in $q_{T}/\Delta p_{T}$. Ultimately, the only way to cure the pathological behavior in the $\Delta p_{T}=0$ limit is to resort to resummation, due to the presence of $x\ln x$ power corrections for $q_{T}/M_{Z}\lesssim x$, as already observed in Frixione:1997ks .
IV.4 Rapidity distributions
Figure 17: Net resummation effect on the inclusive rapidity spectrum, using either the default matching setup of Ref. Ebert:2020dfc with its associated matching uncertainty $\Delta_{\mathrm{match}}$ (left) or the integral-preserving matching setup described in Sec. IV.3 featuring a constrained resummation uncertainty $\Delta_{\mathrm{res}}^{\mathrm{constr.}}$ (right).
We are now in a position to assess the size of resummation effects on the fiducial rapidity distributions studied at fixed order in Sec. III.4. As a baseline, we first assess the impact of the resummation on the inclusive rapidity spectra using the two approaches described in Sec. IV.3 to perform the matching. The results at N3LL$+$NNLO are shown in Fig. 17. For the default matching setup of Ref. Ebert:2020dfc , results from which are shown in the left panel, we find that the integral is compatible with the fixed-order value within the matching uncertainty for all rapidities, providing us with a rapidity-differential version of the check in Eq. (IV.3). The results using the constrained profile scales and the constrained resummation uncertainty introduced in Sec. IV.3 are shown in the right panel. Note that the constraint on the transition points $x_{i}$ was applied at the level of the total inclusive cross section in the ATLAS $Q$ and central $Y$ bin, and not point by point in rapidity. This leads to the finite value of $\Delta_{\mathrm{res}}^{\mathrm{constr.}}/\sigma_{\mathrm{LO}}\approx 0.4\%$ at $2.51.5\leq|y_{Z}|<$ 2.5, where the effect of the beam function $\mu_{B}$ variation changes due to the changing behavior of the underlying PDFs. While it would in principle be possible to apply the constraint point by point in $y_{Z}$ (and $Q$), developing the required automated numerical framework, e.g. along the lines of Ref. Bertolini:2017eui , is beyond the scope of this exercise here. Here we simply point out that for all $y_{Z}$, the integral of the constrained matched prediction is indeed compatible with the fixed-order result within the residual $\Delta_{\mathrm{res}}^{\mathrm{constr.}}$.
Turning to the fiducial $Z$ rapidity spectrum, we first show the effect of resummation for representative values of $\Delta p_{T}$ as a function of $y_{Z}$ in Fig. 18. Our results for $\Delta p_{T}=0$ confirm the observations of Ref. Amoroso:2022lxw : the resummation has a small negative net effect of $-0.4\%$ at central rapidities, then starts to rise at the transition point $|y_{Z}|\sim 1.2$ identified in Sec. III.4, and eventually contributes a small positive net effect of $+0.2\%$ at $|y_{Z}|=2.5$. (Note that in order to read this behavior off from the respective figure in Ref. Amoroso:2022lxw , one must compare the resummation to the unbiased fixed-order result using a recoil prescription.) The fiducial rapidity spectra for $\Delta p_{T}=\pm 5\,\mathrm{GeV}$ feature a very similar trend. We note, however, that the resummation effect at this order is still compatible with zero for most values of $\Delta p_{T}$ and $y_{Z}$ within $\Delta_{\mathrm{res}}^{\mathrm{constr.}}$. To further corroborate these findings, we also compute the net resummation effect on the fiducial $y_{Z}$ spectrum for different $\Delta p_{T}$ using the RadISH resummation code Re:2021con and the settings of Ref. Chen:2022cgv . We find excellent agreement with Fig. 18. This result complements and extends the fixed-order analysis performed in Sect. III.4, showing that when NNLO is used as the baseline, the resummation of linear power corrections has a small effect. This suggests that beyond NNLO, the choice of cuts has an effect at the subpercent level, independently of the presence or absence of linear power corrections, and is likely smaller than the N3LO correction, which has been found to be around $-2\%$ for different choices of cuts in Chen:2022cgv .
In Fig. 19 we show the effect of the resummation on the fiducial rapidity spectrum in the two representative bins $0.4\leq|y_{Z}|\leq 0.6$ and $1.4\leq|y_{Z}|\leq 1.6$ considered earlier in Sec. III.4. In this case we also show results from the default matching setup for comparison. We find that in the more forward bin (gray), which lies past the transition point at $|y_{Z}|\approx 1.2$, the asymmetric and product cuts both show a very similar behavior as a function of $\Delta p_{T}$, with a net resummation effect of $-0.2\%$ (assuming the constrained matching setup) that is compatible with zero within $\Delta_{\mathrm{res}}^{\mathrm{constr.}}$. This is expected since for these values of $|y_{Z}|$, linear power corrections are absent in either case because the lepton pseudorapidity cuts dominate. The picture is different for the central rapidity bin (orange), where the net resummation effect closely resembles that of the total cross section shown earlier in Figs. 15 and 16. In this case, albeit being tiny, the net resummation effect of $\approx-0.3\%$ in the central $|y_{Z}|$ bin is not compatible with zero within uncertainties for both sets of cuts in the region of $\Delta p_{T}$ between $-5$ and $-2\,\mathrm{GeV}$.
Figure 18: Net resummation effect on the fiducial $Z$ rapidity spectrum using asymmetric cuts (left) or product cuts (right) as a function of $|y_{Z}|$ for representative values of $\Delta p_{T}$. For clarity, we only show results for the constrained matching setup. (The results from the default matching setup are not very instructive at this resolution.)
Figure 19: Net resummation effect on the fiducial $Z$ rapidity spectrum using asymmetric cuts (top) or product cuts (bottom) as a function of $\Delta p_{T}$ for two representative bins in $|y_{Z}|$. We show results using the default matching setup (left) and the integral-constrained matching setup (right).
@@ VI Conclusions and Discussion
We also considered product cuts Salam:2021tbm , which replace the separate linear asymmetric cuts on the leading and sub-leading lepton transverse momenta with cuts on their product and the sub-leading lepton instead. For values of $\Delta p_{T}\sim 10$ GeV these cuts mitigate the ambiguities of the perturbative series present with (a)symmetric cuts, which can be particularly relevant for processes with Casimir enhancement, notably Higgs production Billis:2021ecs . As expected, we find that their introduction does not address the pathological behavior for $\Delta p_{T}\rightarrow 0$, which can only be addressed by resummation. We found that resummation indeed removes this effect, although a kink in the cross section is still present for $\Delta p_{T}=0$, essentially due to the way we defined the observable for this study. At NNLO the differences induced by resummation are small for all the values of $\Delta p_{T}$ considered, ranging from 00 to a few GeV, shifting the fixed-order result by sub-percent values that are well below the current experimental uncertainties. An analogous consideration holds also at the level of the rapidity spectra, in agreement with the findings of Ref. Amoroso:2022lxw . It remains to be seen whether such conclusion still holds beyond NNLO, although recent studies Salam:2021tbm ; Chen:2022cgv indicate that the effect of resummation of linear power corrections remains moderate also at N3LO. Given the excellent job done by the available fixed-order codes in reproducing the results of resummation and the substantial effort needed to implement the relevant modifications when modeling signal and background in experimental analyses, we foresee only a marginal improvement in adopting a different set of cuts in the Drell-Yan case for the measurement of fiducial cross sections.
Note added:
The NNLOJET code has been published now in Ref. NNLOJET:2025rno and is available from https://nnlojet.hepforge.org/.